QUESTION 6


A manufacturing company has purchased three assets:

Item
Lathe

Truck

Building
    
Initial cost
$43,000
$25,000
$900,000
Book life
12 years
200,000 miles
50 years
MACRS class
7 years
5 years
39 years
Salvage value
$3,000
$2,000
$100,000
Book depreciation
DDB
Unit production (UP)
SL
 
The truck was depreciated by the units-of-production method.  Usage of the truck was 22,000 miles, 30,000 and 45,000 miles during the first three years, respectively.
 
Calculate the depreciation stated for each asset for the first three years. For the Lathe use DDB.  For the building consider it was bought on May of Year 1.  
Answer #1

1.) Depreciation expense for Truck using Units of production :-

Depreciation rate per miles = (25,000 - 2,000)/200,000 = $0.115

First year Second year Third year
Miles used by Truck (a) 22,000 30,000 45,000
Rate per mile (b) 0.115 0.115 0.115
Depreciation expense (a x b) $2,530 $3,450 5,175

  

2.) Depreciation expense for Lathe using DDB :-

Depreciation rate = (1/12) x 2 = 16.67%

Year 1 depreciation = 43,000 x 16.67% = $7,168.10

Year 2 depreciation = (43,000 - 7,168.10) x 16.67% = $5,973.18

Year 3 depreciation = (43,000 - 7,168.10 - 5,973.18) x 16.67% = $4,977.45

3.) Depreciation expense for Building using SLM :-

Annual depreciation expense = (900,000 - 100,000)/50 = $16,000

Year 1 depreciation = 16,000 x 8/12 = $10,666.67

Year 2 depreciation = $16,000

Year 3 depreciation = $16,000